# Neutrino yield and neutron shielding calculations for a high-power target installed in an underground setting

### v1

@@ Abstract
With the ever increasing beam power at particle accelerator-based facilities for nuclear and particle physics, radioactive isotope production, and nuclear engineering, targets that can withstand this power, and shielding of secondary particles are becoming increasingly important. Here we present Monte Carlo (MC) calculations using the well-established Geant4 software to optimise and predict the antineutrino yield of a 8Li Decay-At-Rest (DAR) source. The source relies on 600 kW of beam power from a continuous wave proton beam impinging on a beryllium target, where spallation neutrons capture on 7Li to produce the 8Li. We further present an in-depth treatment of the neutron shielding surrounding this target. We show that we can produce the high antineutrino flux needed for the discovery-level experiment IsoDAR, searching for “sterile” neutrinos (predicted new fundamental particles) and other beyond standard model physics, while maintaining a neutron flux in the detector that is below natural backgrounds. The methods presented in this paper are easily transferable to other high-power targets and their associated shielding.
I Introduction
@@
Underground laboratories provide areas of exceedingly low background for very sensitive experiments. Placing these experiments in caverns with typically a thousand meters of rock overburden reduces the cosmic muon flux by a factor of several million compared to the rate at the surface. An experiment such as IsoDAR [1, 2], which is designed to place an intense source of neutrinos in close proximity to a large liquid scintillation detector, produces locally extremely high levels of radiation and neutrons, a necessary consequence of the reactions needed to produce the desired neutrino flux.
@@
This poses the question whether it is possible to deploy such a neutrino production facility in an underground laboratory while maintaining the quiet environment for other experiments in the same facility. The answer is yes, but it must be done with careful attention to shielding, regulatory requirements, and containment of all sources of radiation and neutrons.
@@
Strategies for achieving this have been presented in several preceding papers [3, 4, 5], which addressed the deployment of IsoDAR at the Kamioka Observatory, in close proximity to the KamLAND liquid scintillation detector. The present work focuses on a new site, Yemilab, in South Korea, and presents designs appropriate for this site and the planned LSC 2.4 kiloton liquid scintillator detector. More importantly, this paper addresses considerations and general techniques that should be employed in matching such a high-radiation source with the clean and low-background underground environment.
@@
The IsoDAR experiment will use a 60 MeV, 10 mA proton beam, produced by a novel compact cyclotron [6, 7], directed onto a target to produce neutrons, which enter a sleeve of to produce . The rapidly decays ( s) to provide a source of electron antineutrinos ():
@@
The can be detected through inverse beta decay (IBD), , in a large liquid scintillator detector (e.g., the LSC at Yemilab [8]) enabling precision studies of neutrino oscillations, that will definitively elucidate the hints (from, e.g., radioactive-source-based [9, 10, 11] and accelerator-based short baseline neutrino experiments [12, 13]) possibly indicating the existence of “sterile” neutrinos, and allow a range of other beyond-standard-model (BSM) physics searches [1, 14, 15].
@@
The proposed design calls for five years of running, giving a data sample of approximately 1.67 million selected IBD events. This requires the production of antineutrinos/second. Optimising this number requires careful consideration of the target composition and geometry, and this study is the first topic of this paper. The second topic is neutron shielding; the flux within the target is, by design, extremely high and shielding is required to screen the surroundings, both for reasons of safety and to avoid backgrounds in the LSC where a MeV neutron can produce a signal similar to an IBD event. These topics were considered in two previous publications[5, 4]. An update is necessary for several reasons.
@@
The Geant4 ‘Shielding’ package has become available [16], enabling many of the relevant simulations to be done more quickly and accurately; for lower neutron energies the Shielding package uses the high-precision (HP) neutron package which includes the available evaluated neutron data for neutron interactions below 20 MeV. For energies above 20 MeV the Bertini cascade inelastic scattering model [17] is used for energies up to 3 GeV along with the JENDL HE-2007 neutron cross sections available at these higher energies; a direct Geant4-MCNP shielding comparison study has already been done and it has shown very good agreements between the two codes [18]; the Geant4 Shielding package is being used in all the shielding aspects of accelerator, target and irradiation facilities workshops [19] since 2002 as part of the ’Intercomparison of Medium-energy Neutron Attenuation in Iron and Concrete’ project [20].
@@
The previous accounts considered only the KamLAND site; here, we address Yemilab-specific issues, however the methods employed can be applied to the requirements of any specific site.
5.
@@
The general layout of the experiment is shown in Figure 1. The beam from the cyclotron (off the figure to the right, in an existing cavern specifically constructed for it by the Yemilab Laboratory) is bent through before striking the target. This geometry was chosen to give space to enable the target assembly to be withdrawn and replaced, which we anticipate will be necessary several times during the lifetime of the experiment. It has the additional benefit of reducing the background from high energy neutrons directed towards the detector, as will be discussed later (see Fig. 9). The LSC is shielded from the target by steel and concrete: the thickness of these layers is an important design choice, as increasing the thickness reduces the neutron background but also the antineutrino signal (since the antineutrino yield is isotropic, increasing shielding increases the distance from the detector, and the flux drops as 1/r2).
@@
The IsoDAR target design presented in Ref. [5] and updated in Ref. [3] has been redesigned, as shown in Figure 2.
@@
At the vacuum interface, the proton beam impinges on a 3 mm thick hemispherical beryllium shell. Behind it lie two more concentric, 3 mm thick, beryllium shells separated by 2 mm deep channels of heavy water. Finally, the centre is comprised of a solid beryllium block. Full details, including strategies to avoid stopping of protons in beryllium layers of the target, which is known to lead to embrittlement because of poor solubility of hydrogen in beryllium [21], are given in an upcoming publication on the mechanical design and fluid dynamics of the target and cooling.
II Methods
To simulate all stages of the process we use the Geant4 library[22]. Originally developed for the description of particle physics detectors, it has since been expanded and applied to many fields. We use version Geant4-11.0.2. The shielding physics list was used for the shielding calculations, and the binary cascade and QGSP options are used for the reactions relevant to antineutrino production. The “high precision” cross section libraries for neutron interactions are included, as these processes turn out to have an important contribution.
@@ II Methods
In order to maximise the production of antineutrinos we must maximise that of the precursor neutrons. The disintegration of was the method by which the neutron was first seen[23] and is still the most efficient way of producing neutrons using a small accelerator. (Lithium also gives neutrons, but is structurally unworkable, with a low melting point.)
@@
The target constitutes a very high intensity neutron source, and shielding must be put in place to attenuate these penetrating particles beyond the production sleeve. The IsoDAR experiment will be deep underground and will be controlled remotely, so exposure of personnel is not an issue. Components will become activated, but maintenance work will be possible after a suitable cooling-off period. The two principal concerns are for the activation in the surrounding rock, which must remain below a statutorily-required level (discussed later) on completion of the experiment, and backgrounds in the LSC detector, where a high energy neutron interaction, with possible subsequent capture, can mimic the signature of an antineutrino event.
@@
Iron, which has a high interaction cross section for high energy neutrons (above MeV).
@@
A typical configuration considered was a thick layer of iron, which slowed down the fast neutrons, followed by a layer of boron loaded concrete to remove the resulting slow neutrons. Gadolinium loaded polyethylene was also considered since it has properties similar to Boron loaded concrete.
Considering the background in the detector, low energy neutrons below MeV are not a problem as they do not resemble the signal from an antineutrino-induced IBD or elastic scattering event, so the second shielding layer is not necessary for this purpose.
III Results
III-A Neutron and antineutrino yield
The new IsoDAR design features a hemispherical beryllium target, as described in Section I.
@@ III-A Neutron and antineutrino yield
The previous design[5] used a cylindrical sleeve, but simulations showed that the neutron production was, as one would expect, roughly spherical, and so the amount of expensive sleeve material can be reduced by removing the least productive volumes. The optimised radius is 75 cm and the length 110 cm, so it is an ellipsoid rather than a sphere (though for some simulations a sphere is used for simplicity).
@@
The previous design studies [5] showed that the optimum sleeve material was a mixture of beryllium and isotopically pure , with a Be mass fraction of 75%. It is formed of 1 mm diameter beryllium spheres in lithium. A FLiBe eutectic mix was also considered, but is less performant [5].
@@
The previous design achieved a production rate of 0.018 per incident proton. The new nested-shell target design, necessary for cooling, initially reduced this by a factor . This figure, which is a key measure of performance, has been increased again by optimisation of the target and sleeve geometry, and by increasing the isotopic purity of the lithium. 6Li has a very large neutron cross section, and even small amounts will have a serious negative effect. Natural lithium is 92.5% pure 7Li, and we need this to be increased to well over 99%, as shown in Figure 4. A purity of 99.99% will give 0.016 per proton, which is sufficient for the experiment, and 99.995% will give 0.020. A higher purity would perform even better, but we take the lower number as our baseline, due to cost and availability.
III-B Shielding

### v2

@@ Abstract
With the ever increasing beam power at particle accelerator-based facilities for nuclear and particle physics, radioactive isotope production, and nuclear engineering, it becomes increasingly important to have targets that can withstand this power, and shielding to block the secondary particles produced. Here we present Monte Carlo (MC) calculations using the well-established Geant4 software to predict the antineutrino yield of a 8Li Decay-At-Rest (DAR) source. The source relies on 600 kW of beam power from a continuous wave proton beam impinging on a beryllium target, where spallation neutrons are captured by 7Li to produce the 8Li. We further present an in-depth treatment of the neutron shielding surrounding this target. We show that we can produce the high antineutrino flux needed for the discovery-level experiment IsoDAR, searching for “sterile” neutrinos (predicted new fundamental particles) and other beyond standard model physics, while maintaining a neutron flux in the detector that is below natural backgrounds. The methods presented in this paper are easily transferable to other high-power targets and their associated shielding.
1 Introduction
@@
Underground laboratories provide areas of exceedingly low background for very sensitive experiments [Formaggio:Background]. Placing these experiments in caverns with typically a thousand meters of rock overburden reduces the cosmic muon flux by a factor of several million compared to the rate at the surface. For experiments that cannot make use of a very narrow beam time window to minimize the effects of atmospheric muons, this shielding is absolutely necessary. An example of such an experiment is IsoDAR [alonsoNeutrinoPhysicsOpportunities2022, winklehnerIsoDARYemilabDefinitive2023], designed to study neutrinos and search for rare phenomena. The configuration of this experiment places an intense neutrino source in close proximity to a large liquid scintillation detector. Extremely high fluxes of neutrons and other radiation are produced in the vicinity of the target, a necessary consequence of the reactions needed to produce the desired neutrino flux.
@@
This poses the question whether it is possible to deploy such a neutrino production facility in an underground laboratory while maintaining the quiet environment for other experiments in the same facility. The answer is yes, but it must be done with careful attention to shielding, regulatory requirements, and containment of all sources of radiation and neutrons. As explained later, this implies fluxes that produce rock activation below 10 Bq/g (for Korea, or 0.1 Bq/g for Japan), and very small numbers of neutron-induced background events in the detector.
@@
Strategies for achieving this have been presented in several previous papers [alonsoIsoDARYemilabReport2022, Shieldingpaper, Bungau_2019], which addressed the deployment of IsoDAR at the Kamioka Observatory, in close proximity to the KamLAND liquid scintillation detector. The present work focuses on a new site, Yemilab, in South Korea, and presents designs appropriate for this site and the planned EYE 2.4 kiloton liquid scintillator detector (formerly known as the Liquid Scintillator Counter, or LSC). More importantly, this paper addresses considerations and general techniques that should be employed in matching such a high-radiation source with the clean and low-background underground environment.
@@
The IsoDAR experiment will use a 60 MeV, 10 mA proton beam, produced by a novel compact cyclotron [winklehnerNewFamilyHighcurrent2024, winklehnerOrderofmagnitudeBeamCurrent2022], directed onto a target to produce neutrons, through reactions such as beryllium break-up. These neutrons enter a surrounding sleeve of beryllium, for additional neutron production, and . Neutron capture on results in , which rapidly decays ( s) to provide a source of electron antineutrinos ():
@@
The can be detected through inverse beta decay (IBD), , in a large liquid scintillator detector (e.g. the EYE at Yemilab [Seo:2023xku]) enabling a sensitive search for neutrino oscillations at short distances, which will definitively elucidate the hints (from, e.g., radioactive-source-based [gallex, sage, Barinov:2022wfh] and accelerator-based short baseline neutrino experiments [lsnd3, MiniBooNENeutrino2021]) possibly indicating the existence of “sterile” neutrinos, and allow a range of other beyond-standard-model (BSM) physics searches [alonsoNeutrinoPhysicsOpportunities2022, Waites:2022tov, Hostert:2022ntu].
@@
The proposed design calls for five years of running, giving a data sample of approximately 1.67 million selected IBD events. This requires the production of antineutrinos/second. Optimising this number required careful consideration of the target composition and geometry, described in [Bungau_2019] and references therein, and the study of the neutron and antineutrino yield is the first topic of this paper. The second topic is neutron shielding; the flux within the target is, by design, extremely high and shielding is required to screen the surroundings, both for reasons of safety and to manage backgrounds in the EYE.
An MeV requirement on candidate single-flash-signal events in the fiducial volume, which starts around 3 m from the outer wall of the detector, is planned. Around MeV, the background behavior changes[Formaggio:Background, Abe, Suekane]. Below 3 MeV, there are large number of backgrounds from radioactive decays in the PMTs and the inner detector tank, such as and . On the other hand, contributions from radioactivity are small above 3 MeV (see Ref. [alonsoNeutrinoPhysicsOpportunities2022] for a detailed analysis). This requirement was motivated by the goal of reducing backgrounds from natural radioactivity to an acceptable level, and still allows 0.07 Hz of inverse beta decay interactions and 0.2 mHz of elastic scattering events to be accepted for physics studies. Although introduced for other reasons, this cut has the benefit of also eliminating neutrons from the target that penetrate the outer regions of the detector, but then deposit less than 3 MeV of energy—hence this discussion focuses on neutrons above 3 MeV, aiming for a rate of Hz given the shielding we describe, which is three orders of magnitude below the rate for elastic scatters, making this background negligible for neutrino analyses and allowing high sensitivity for Beyond Standard Model event searches.
Both optimising antineutrino production and shielding have been considered in two previous publications[Bungau_2019, Shieldingpaper]. An update is necessary for several reasons.
@@
The previous accounts considered the KamLAND site, and here we address requirements for Yemilab. ( However the methods employed can be applied to any site, depending on physical and regulatory requirements.)
@@
The general layout of the experiment is shown in Figure 1. The beam from the cyclotron (off the figure and about 20 meters to the right, in an existing cavern specifically constructed for it by the Yemilab Laboratory) is bent through before striking the target. The target assembly, beryllium hemispheres and cooling circuits shown in Figure 2 – referred to as the "torpedo" – is inserted through the downstream end of the beam pipe. The outer beryllium hemisphere is 20 cm diameter, the beam is spread out to distribute the beam power over the entire target surface. This maintains the maximum power density at less than 8 kW/cm2, a safe limit for cooling design. The cart with pumps and heat exchangers is seen behind the shielding in Figure 1. This geometry has the advantage of providing space to enable the target assembly to be withdrawn and replaced from the downstream end of the beam pipe with remote handlers, which we anticipate will be necessary several times during the lifetime of the experiment. Boreholes drilled into the cavern walls serve as storage for new and spent targets. As the proton beam is thus directed away from the EYE it has the additional benefit of reducing the background from high energy neutrons directed towards the detector, as will be discussed later (see Fig. 12). The EYE is shielded from the target by steel and concrete: the thickness of these layers is an important design choice, as increasing the thickness reduces the neutron background but also the antineutrino signal (since the antineutrino yield is isotropic, increasing shielding increases the distance from the detector, and the flux drops as 1/r2). In the present design the center of the fiducial volume of the EYE is 17.6 meters from the target, including a 4-meter slab of steel and the beamline magnet bringing the beam into the target.
@@
The IsoDAR target design presented in Ref. [Bungau_2019] and updated in Ref. [alonsoIsoDARYemilabReport2022] has been redesigned, as shown in Figure 2.
@@
At the vacuum interface, the proton beam impinges on a 3 mm thick hemispherical beryllium shell. Behind it lie two more concentric, 3 mm thick, beryllium shells separated by 7 mm deep channels of heavy water, pumped and cooled by a remote external circulator. Finally, the centre is comprised of a solid beryllium block. The design contains 480 kg of and 1437 kg of Be in the sleeve, another 44.5 kg of Be in the target, and 18.65 kg of heavy water. Full details, including strategies to avoid stopping of protons in beryllium layers of the target, which is known to lead to embrittlement because of poor solubility of hydrogen in beryllium [Rinckel_2012], will be given in a projected future publication on the mechanical design and fluid dynamics of the target and cooling.
The cylindrical Li/Be sleeve of the previous design has been replaced by a (roughly) spherical section, as shown in Figure 3, following the optimisation analysis developed in [Bungau_2019].
2 Methods
To simulate all stages of the process we use the Geant4 code[agostinelli, ALLISON2006, ALLISON2016]. Originally developed for the description of particle physics detectors, it has since been expanded and applied to many fields. We use version Geant4-11.0.2.
Different stages of the simulation require different Geant4 physics lists, as no single option is suitable for all processes involved. For the initial stage of neutron and production, we employed QGSP_BIC_AllHP, which combines the binary cascade model for higher energies with high-precision, data-driven models (ParticleHP) for low-energy proton and neutron interactions based on ENDF and TALYS cross-section libraries. This physics list is recommended for applications such as BNCT (Boron Neutron Capture Therapy) and provides reliable agreement with experimental data.
For the shielding studies, we used the Shielding physics list [Shieldinglist]. Originally developed for neutron penetration and ion–ion collision studies, it is also applicable to calorimetry and low-background experiments. Its high-energy treatment derives from FTFP_BERT, with radioactive decay processes included to account for background radiation, while retaining the same evaluated neutron cross-section libraries as the high-precision models. For energies above 20 MeV, the Bertini cascade inelastic scattering model [Wright] is used for energies up to 3 GeV with parameters tuned with the help of the JENDL HE-2007 neutron cross sections [JENDL] available at these higher energies.
@@ 2 Methods
In order to optimise the production of antineutrinos we must maximise that of the precursor neutrons. The disintegration of 9Be was the method by which the neutron was first seen[Chadwick] and is still the most efficient way of producing neutrons using a small accelerator. (Lithium also gives neutrons, but is structurally unworkable, with a low melting point.)
@@
The target constitutes a very high intensity neutron source, and shielding must be put in place to attenuate these penetrating particles beyond the production sleeve. The IsoDAR experiment will be deep underground and will be controlled remotely, so exposure of personnel is not an issue. Components will become activated, but maintenance work will be possible after a suitable cooling-off period. The two principal concerns are for the activation in the surrounding rock, which must remain below a statutorily-required level (discussed later) on completion of the experiment, and backgrounds in the EYE detector, where a high energy neutron interaction, with possible subsequent capture, can mimic the double-coincidence signature of an antineutrino event.
@@
Iron, which has a large interaction cross section for high energy neutrons (above MeV).
@@
A typical configuration considered was a thick layer of iron, which softened the spectrum, followed by a layer of boron loaded concrete to remove the resulting slow neutrons.
For environmental shielding inside the cavern and mitigation of rock activation, attenuation of all neutrons, respective of energy, is important. For this, it is optimal to provide a combination of iron to efficiently reduce the flux of higher-energy neutrons through inelastic collisions, and absorption of slower neutrons through boron loaded concrete. For backgrounds in we only need consider neutrons above 3 MeV, as discussed in Section 3.2.5, so the concrete attenuation is not really required. But in addition, the buffer and veto layers in the detector provide ample attenuation of neutrons emerging from the large steel block between the target and the detector.
2.1 Validation
Although Geant4 is very widely used for simulation in particle physics experiments, its predictions need to be validated and uncertainties estimated, and the impact of such uncertainties on the proposed experiment evaluated.
2.1.1 Validation of neutron production rates
We have conducted a comprehensive validation of the Geant4 physics models relevant to low-energy proton interactions with beryllium using experimental data from Osipenko et al.[Osipenko] at 62 MeV and from Tilquin et al. [tilquin] at 23-80 MeV, which show good agreement between Geant4 predictions and measured neutron yields and have presented these in previous publications [Shieldingpaper, Bungau_2019]. We demonstrated that, with appropriate physics models selection, Geant4 accurately reproduces neutron yields for proton energies between 23 MeV and 55 MeV, with only a slight underestimation at higher energies. In Fig. 4 (not previously pubished) we compare predicted neutron yields from thick Be targets with experimental data from [tilquin].
The good agreement observed provides strong support for using Geant4 to model neutron production in the energy range relevant to our experiment.
If the Geant4 predicted yields do turn our to be optimistic, by 10% or even higher (though we do not expect that), the impact on the experiment is not catastrophic. Lower rates will mean that we need to run the experiment for a longer time or, alternatively, to publish results with somewhat larger statistical uncertainties.
2.1.2 Validation of shielding studies
The Geant4 ‘Shielding’ physics list [Shieldinglist] has been continuously used in all the shielding aspects of accelerator, target and irradiation facilities workshops [Hirayama8, Matsuda, Conf1, NEA1, NEA2, NEA3] since 2010 as part of the ‘Intercomparison of Medium-energy Neutron Attenuation in Iron and Concrete’ project [Hirayama6]. These studies involved an intercomparison of the available Monte Carlo codes which can be used in shielding calculations, and they show how the Geant4 Shielding physics list has evolved and improved over the years since it was added to the comparison list in 2010 [Shieldinglist]. It has been used not only for a direct comparison against other Monte Carlo codes predictions, but also for validation studies in direct comparisons with experimental data, showing very good agreements [NEA2]. For lower neutron energies, the Shielding physics list uses the high-precision (HP) neutron physics model which includes the available evaluated neutron data for neutron interactions below 20 MeV.
These studies are cited in the present work to clarify that the validation of the relevant Geant4 models has already been completed and demonstrated to be reliable within the energy range of interest.
3 Results
3.1 Neutron and antineutrino yield
The new IsoDAR design features a hemispherical beryllium target, as described in Section 1.
@@ 3.1 Neutron and antineutrino yield
The previous design [Bungau_2019] used a cylindrical sleeve, but simulations showed that the neutron production was, as one would expect, roughly spherical, and so the amount of expensive sleeve material can be reduced by removing the least productive volumes. An optimization study was performed to determine the best shape and size of the target sleeve to maximise production. The study used a three-dimensional mesh to quantify the spatial distribution of within the volume of the sleeves. By calculating the content in each mesh cell, regions that contribute minimally to the overall yield were identified. Mesh cells with concentrations below a predefined threshold were systematically excluded, allowing effective sleeve geometry to be inferred. This approach enables the design of a sleeve that maximizes production while minimizing unnecessary material, providing a clear guideline for the target geometry in experimental setups. The results are shown in Figure 5.
The optimised design can be approximated by a spherical shape of 60 cm radius, excluding the outer stainless steel shell.
@@
The previous design studies [Bungau_2019] showed that the optimum sleeve material was a mixture of beryllium and isotopically pure , with a Be mass fraction of 75%. It is formed of 1 mm diameter beryllium spheres in lithium. A FLiBe eutectic mix was also considered, but is less performant [Bungau_2019].
@@
The previous design achieved a production rate of 0.018 per incident proton. The new nested-shell target design, necessary for cooling, initially reduced this by a factor . This figure, which is a key measure of performance, has been increased again by optimisation of the target and sleeve geometry. We note that 6Li has a very large neutron cross section, and even small amounts will have a serious negative effect. Natural lithium is 95.15% pure 7Li, and we need this to be increased to well over 99%, as shown in Figure 6. A purity of 99.99% will give 0.016 per proton, which is sufficient for the experiment, and 99.995% will give 0.020. Isotopically purified lithium is commercially available due to its importance in molten salt reactors and in the fusion program. A higher purity would perform even better, but we take the lower number as our baseline, due to cost and availability.
is an ideal single-isotoped neutrino source as the decay produces antineutrinos at very high energy compared to other beta decays. The IBD process in the liquid scintillator essentially eliminates backgrounds from neutrons, neutrinos and cosmic rays by requiring two signals in coincidence, from the first flash due to the positron and subsequent second flash from the neutron capture. The IBD reaction requires a neutrino energy of 1.8 MeV, quite high for a beta decay. The spectrum is peaked at 6 MeV whereas the other isotopes produced (in smaller quantities) in the target produce much lower energies. Anti-neutrinos from tritium decay only go up to 0.019 MeV, from up to 0.55 MeV. There will be a small contribution from , which goes up to 3.5 MeV and is produced in very small amounts but otherwise the antineutrino flux above 2 MeV will only come from decay.
Also the spectrum is well understood: it has been extensively studied due to its role in solar neutrino physics, so the systematic uncertainty from the theoretical prediction of the spectrum is very low. It is widely used for tests for BSM physics because the shape of the spectrum is so well known. See for example[spectrum1, spectrum2] which describe a test for tensor currents and discusses the SM prediction.
3.2 Shielding
Our shielding assessments are made with several goals in mind, each one having its own set of requirements.
1.
Biological shielding, to enable personnel access as quickly as possible after beam shutoff. While the beam is running radiation levels will be too high for personnel access, regardless of shielding design. But it is important to minimize short-lived activity of items in the IsoDAR caverns, considered in the sections 3.2.2 and 3.2.4 on component activation and tritium production.
2.
Rock activation, as neutrons that penetrate the rock surfaces can cause activation, and some long-lived isotopes, such as may become a problem if activity persists beyond the lifetime and decommissioning of the experiment (Section 3.2.3.)
- 3.
In the studies for siting IsoDAR at the Kamioka Observatory, the high sodium content of the Inishi rock at this site dominated the shielding requirements [Shieldingpaper], and led to the very large block we will describe.